# Task-conditioned adaptation of visual features in multi-task policy learning

### v1

@@ 4 Experiments
Visualizing the influence of task-conditioned adaptation — Sequences of visual frames used in this experiment are taken from a held-out set of expert trajectories not used at training time. We visualize here the attention map of the last layer of the vision encoder. To this end, we sum attention maps for all tokens and all heads, and normalize them between 00 and $1$. These visualizations are shown in Figures 9. The first row shows a sequence of visual frames and below, for each model variant (No Adapter, Middle Adapter (NC), Middle Adapter (C)), one can see the attention map overlaid on top of the visual frame and displayed below as a colored heatmap. As can be seen, the middle adapters help focus the attention on the most important parts of the image compared with vanilla VC-1 attention without adapters. When adapters are not conditioned (NC), they tend to produce very narrow attention. Conditioning on the task at hand keeps the focus on important regions and leads to covering the entire objects of interest and important agent parts. Most importantly, Figure 9 shows that, when adapters are conditioned on the task embedding, more attention is put on the final goal in all frames, while this is not the case for unconditioned adapters. Another visualization is shown in Figure 11 in the appendix.
@@ Appendix B Additional ablation studies
Table 2: Known task  — Additional ablation studies: Validation and test performance on known tasks of different neural variants. Row (a) is equivalent to row (f) in Table 1 of the main paper. When using conditioned adapters, giving the task embedding as input to the policy is not necessary. Our introduced tokens aggregation layer is better than using the representation of the ’CLS’ token. Finally, when training a conditioned top adapter, middle adapters are still important and bring a boost in performance. Cond $\pi$: policy conditioned on the task embedding – C: Conditioned – ’CLS’ token: using the ’CLS’ token representation as the frame embedding fed to the policy.
@@
Table 3: Known task  — Impact on other visual backbones: Validation and test performance on known tasks for two additional visual backbones (PVR [23] and MVP [27]). Our task-conditioned adapters improve the extracted visual features in both cases, leading to higher multi-task policy performance.
@@
Table 5: Known task  — Diversity of known tasks: Validation and test performance on MetaWorld known tasks when our approach with task-conditioned adapters is either trained on the three considered benchmarks (Adroit, DMC, MetaWorld), or on tasks from MetaWorld only. As expected, the model trained on known tasks from MetaWorld only reaches higher performance.
@@
Table 6: Few-shot  — Performance of a finetuned baseline (Ft.) and task embedding search for a policy either trained on MetaWorld only or all 3 benchmarks. $t_{i}^{u}$ refers to the $i$-th unknown task.
@@ Appendix D Visualizing the influence of task-conditioned adaptation
Figure 12: Task-related information inside visual embeddings: t-SNE plots of visual embeddings for a set of frames for DMC Stand and Walk tasks. We chose these tasks for their visual similarity, making it very hard to distinguish between them from vision only. Conditioning of the middle adapters leads to two properly separated clusters, showing the insertion of task-related information into the visual embedding.
Figure 12: Task-related information inside visual embeddings: t-SNE plots of visual embeddings for a set of frames for DMC Stand and Walk tasks. We chose these tasks for their visual similarity, making it very hard to distinguish between them from vision only. Conditioning of the middle adapters leads to two properly separated clusters, showing the insertion of task-related information into the visual embedding.
@@ Appendix F Diversity of known tasks
Tab. 6 (c) and Tab. 5 (b) shows a model trained on MetaWorld only, which performs better on MetaWorld than models trained on all 3 benchmarks (the domain gap between them is large). The lower performance on MetaWorld when training on all 3 benchmarks is largely outweighed by the ability to address Adroit and DMC.
@@ Appendix H Architecture details
Middle adapters — one adaptation module $\alpha_{i}$ is inserted after each self-attention layer inside $\phi$. It is composed of $2$ fully-connected layers with respectively $384$ and $768$ neurons. A GELU activation function is applied to the output of the first layer. The input to a middle adapter is the concatenation of the task embedding and a token representation from the previous self-attention layer. It thus processes all tokens as a batch.

### v2

@@ 4 Experiments
— Visualizing the influence of task-conditioned adaptation — Sequences of visual frames used in this experiment are taken from a held-out set of expert trajectories not used at training time. We visualize here the attention map of the last layer of the vision encoder. To this end, we sum attention maps for all tokens and all heads, and normalize them between 00 and $1$. These visualizations are shown in Figures 8. The first row shows a sequence of visual frames and below, for each model variant (No Adapter, Middle Adapter (NC), Middle Adapter (C)), one can see the attention map overlaid on top of the visual frame and displayed below as a colored heatmap. As can be seen, the middle adapters help focus the attention on the most important parts of the image compared with vanilla VC-1 attention without adapters. When adapters are not conditioned (NC), they tend to produce very narrow attention. Conditioning on the task at hand keeps the focus on important regions and leads to covering the entire objects of interest and important agent parts. Most importantly, Figure 8 shows that, when adapters are conditioned on the task embedding, more attention is put on the final goal in all frames, while this is not the case for unconditioned adapters. Another visualization is shown in Figure 10 in the appendix.
@@ 5 Conclusion
Acknowledgement — We thank ANR for support through AI-chair grant “Remember” (ANR-20-CHIA-0018).
@@ Appendix B Additional ablation studies
Table 2: Known task  — Additional ablation studies: Validation and test performance on known tasks of different neural variants. Row (a) is equivalent to row (f) in Table 1 of the main paper. When using conditioned adapters, giving the task embedding as input to the policy is not necessary. Our introduced tokens aggregation layer is better than using the representation of the ’CLS’ token. Finally, when training a conditioned top adapter, middle adapters are still important and bring a boost in performance. Cond $\pi$: policy conditioned on the task embedding – C: Conditioned – ’CLS’ token: using the ’CLS’ token representation as the frame embedding fed to the policy. Performance is reported as mean $\pm$ std over $3$ training runs (seeds).
@@
Table 3: Known task  — Impact on other visual backbones: Validation and test performance on known tasks for two additional visual backbones (PVR [23] and MVP [27]). Our task-conditioned adapters improve the extracted visual features in both cases, leading to higher multi-task policy performance. Performance is reported as mean $\pm$ std over $3$ training runs (seeds).
@@
Table 5: Known task  — Diversity of known tasks: Validation and test performance on MetaWorld known tasks when our approach with task-conditioned adapters is either trained on the three considered benchmarks (Adroit, DMC, MetaWorld), or on tasks from MetaWorld only. As expected, the model trained on known tasks from MetaWorld only reaches higher performance. Performance is reported as mean $\pm$ std over $3$ training runs (seeds).
@@
Table 6: Few-shot  — Performance of a finetuned baseline (Ft.) and task embedding search (TE opt.) for a policy either trained on MetaWorld only (MV) or all 3 benchmarks (All 3). $t_{i}^{u}$ refers to the $i$-th unknown task. Performance is reported as mean $\pm$ std over $3$ training runs (seeds).
@@ Appendix D Visualizing the influence of task-conditioned adaptation
Figure 11: Task-related information inside visual embeddings: t-SNE plots of visual embeddings for a set of frames for DMC Stand and Walk tasks. We chose these tasks for their visual similarity, making it very hard to distinguish between them from vision only. Conditioning of the middle adapters leads to two properly separated clusters, showing the insertion of task-related information into the visual embeddings.
@@ Appendix F Diversity of known tasks
Table 6 (c) and Table 5 (b) show a model trained on MetaWorld only, which performs better on MetaWorld than models trained on all 3 benchmarks (the domain gap between them is large). The lower performance on MetaWorld when training on all 3 benchmarks is largely outweighed by the ability to address Adroit and DMC.
@@ Appendix H Architecture details
Middle adapters — one adaptation module $\alpha_{l}$ is inserted after each self-attention layer inside $\phi$. It is composed of $2$ fully-connected layers with respectively $384$ and $768$ neurons. A GELU activation function is applied to the output of the first layer. The input to a middle adapter is the concatenation of the task embedding and a token representation from the previous self-attention layer. It thus processes all tokens as a batch.